OpenStudy (anonymous):

HELP PLEASE......Write the equation in slope-intercept form of the line that has a slope of 2 and y-intercept of -2

OpenStudy (jim766):

y = mx + b m is slope b is y intercept, just substitute in y = 2x - 2 make sense?

OpenStudy (anonymous):

yes thx

OpenStudy (anonymous):

1 more tho...Write the equation in slope-intercept form of the line that has a slope of 3 and contains the point (4, 9).

OpenStudy (jim766):

in this one the give you the slope, but now y intercept, so you have to figure it out. but they do give you a point to use x y (4, 9) y = mx + b we just sub in what the gave us and solve for b 9 = 3(4) + b do you know how to get the b alone?